OpenStudy (anonymous):

I've been stuck on this problem for days. Solve the system by elimination. { -2x + 2y + 3z = 0; -2x - y + z = -3; 2x + 3y + 3z = 5. Any help would be nice! Normally I wouldn't be this stuck :/

OpenStudy (anonymous):

Really need help on this :/

OpenStudy (owlcoffee):

I'll help you out.

OpenStudy (anonymous):

Thank you!

OpenStudy (dumbcow):

key is to eliminate a variable to get it into a system of 2 equations with 2 variables, then solve that smaller system with elimination. then by substitution you can find the 3rd variable @Owlcoffee will demonstrate :)

OpenStudy (owlcoffee):

So we have: -2x + 2y + 3z = 0 -2x - y + z = -3 2x + 3y + 3z = 5 Let's first pick two equations here, i'll pick the first and the second: -2x + 2y + 3z = 0 -2x - y + z = -3 I'll multiply the first one by -1 just to eliminate the x's. ending up with: 2x - 2y - 3z = 0 -2x - y + z = -3 I will sum this equations, so the 2x and the -2x will simplify, i'll end with an equation looking like this: -3y-2z=-3 We have just created an equation with only two variables, but we need to make another one :( I encourage you to fabricate the other ecuation, it's the same idea. :)

OpenStudy (anonymous):

Okay.. I did: -2x - y + z = -3 2x + 3y + 3z = 5 from that I got 2y + 4z = 2. So would I write the two new equations like -3y - 2z = 3 2y + 4z = 2

OpenStudy (owlcoffee):

Good! now solve that 2x2 system :)

OpenStudy (anonymous):

Okay! I got -6 y - 4z = 6 2y + 4z = 5 then -4y = 8 and divided by -4 which got y = -2. So now do I have to find x and z?

OpenStudy (anonymous):

You have a equation ! So if you get y the number you have can find both of them !:)

OpenStudy (owlcoffee):

yes, on the 2nd system, replace the value of y to find z. and on the first 3x3, replace both z and y to get the value of x.

OpenStudy (anonymous):

So on the -6y - 4z =6 I would replace the y? or the -2x - y + z= -3? A little confused on this part :s